### v3

# Applying Non-negative Matrix Factorization with Covariates to the Longitudinal Data as Growth Curve Model

Abstract
Using Non-negative Matrix Factorization (NMF), the observed matrix can be approximated by the product of the basis and coefficient matrices. Moreover, if the coefficient vectors are explained by the covariates for each individual, the coefficient matrix can be written as the product of the parameter matrix and the covariate matrix, and additionally described in the framework of Non-negative Matrix tri-Factorization (tri-NMF) with covariates. Consequently, this is equal to the mean structure of the Growth Curve Model (GCM). The difference is that the basis matrix for GCM is given by the analyst, whereas that for NMF with covariates is unknown and optimized. In this study, we applied NMF with covariance to longitudinal data and compared it with GCM. We have also published an R package that implements this method, and we show how to use it through examples of data analyses including longitudinal measurement, spatiotemporal data and text data. In particular, we demonstrate the usefulness of Gaussian kernel functions as covariates.
keywords:
Growth Curve Model (GCM), Non-negative Matrix Factorization (NMF), Non-negative Matrix tri-Factorization (tri-NMF), Soft Clustering, Topic Model, Varying Coefficients.
1 Introduction
Non-negative Matrix Factorization (NMF) was proposed by Lee and Seung [15] to approximate an observed matrix consisting of non-negative elements as the product of a basis and coefficient matrices. It is often regarded as a dimension-reduction method because it represents multidimensional data with lower-dimensional bases; however, it can also be considered a regression model regarding approximating bases and coefficients. However, because the coefficients are non-negative and can be interpreted as proportions or probabilities, they are used for soft clustering. NMF is a versatile analytical method for both supervised and unsupervised machine learning. NMF has been applied in various fields, such as image data object recognition by Lee and Seung [15], music transcription and audio source separation by Hayashi et al [12], and text data by Berry et al [2] and Mifrah and Benlahmar [16].
Optimizing all elements of the basis and coefficient matrices as parameters is inherently very difficult; however, by using the update formula derived by Lee and Seung [14], the objective function is minimized every time the update formula is applied, provided that the initial value is given. Therefore, it is an effective analysis method for big data, as numerical calculations are less likely to collapse, even when the dimension of the observed vector is high. Moreover, there is an advantage since the analyst can easily devise add constraints and alike as needed.
However, in ordinary NMF only the observation matrix, not the covariates, is used. Therefore, predictions based on the background or risk factors are not possible. Therefore, we consider expressing the coefficient matrix as the product of the parameter matrix and the covariate matrix and approximate the observation matrix by the product of three matrices: the basis matrix, the parameter matrix, and the covariate matrix, in the same manner as the mean structure of the Growth Curve Model (GCM) proposed by Potthoff and Roy [18]. Although the decomposition by three non-negative matrices was also proposed by Ding et al [6] and is called Non-negative Matrix tri-Factorization (tri-NMF), see Gillis [9], we emphasize “covariates” and call it NMF with covariates.
NMF with covariates has the advantage of making predictions. However, using the covariate matrix limits the possible values of the coefficient matrix, resulting in lower approximation accuracy than when covariates are not used. To address this issue, Satoh [20] proposed using a Gaussian kernel function, which is known to have high explanatory power, as a covariate (see Chapter 14 of Murphy [17]) and suppresses the decrease in approximation accuracy.
Section 2 introduces NMF with covariates and discusses the covariates composed of Gaussian kernel functions. Section 3 presents the update formulas for optimization in the tri-NMF framework. In Section 4, we describe its relationship with GCM. Section 5 presents certain applications of NMF with covariates on real data, available on GitHub as an R package (R Core Team [19]).
2 Non-negative Matrix Factorization with covariates
For the observation matrix $Y=({\bm{y}}_{1},\dots,{\bm{y}}_{N})=(y_{p,n})_{P\times N}$ consisting of non-negative $P$ variables from $N$ individuals, we consider an appropriate $Q\leq\min(P,N)$ and approximate it by the product of the basis matrix $X=({\bm{x}}_{1},\dots,{\bm{x}}_{Q})=(x_{p,q})_{P\times Q}$ and the coefficient matrix $B=({\bm{b}}_{1},\dots,{\bm{b}}_{N})=(b_{q,n})_{Q\times N}$, both having non-negative elements:
[…]
This approximation of the observation matrix by the product of the basis matrix and coefficient matrix, or NMF, was proposed by Lee and Seung [15, 14]. Next, we explain the coefficient matrix $B$ using covariates. In other words, we use the known covariate matrix $A=({\bm{a}}_{1},\dots,{\bm{a}}_{N})=(a_{k,n})_{R\times N}$ and the parameter matrix $\Theta=(\theta_{q,k})_{Q\times R}$, both having nonnegative elements, to write
[…]
Thus, $Y\approx XB=X\Theta A$. Additionally, for individual $n$, we have
${\bm{y}}_{n}\approx X{\bm{b}}_{n}=X\Theta{\bm{a}}_{n},\qquad n=1,\ldots,N,$ (3)
This approximation model, expressed as the product of three matrices, is included in the tri-NMF, proposed by Ding et al [6]. Note that in tri-NMF, the three matrices are generally unknown; however, the covariate matrix in this study is known. Thus, for a new covariate ${\bm{a}}$, we obtain the coefficient vector ${\bm{b}}=\Theta{\bm{a}}$ and predict the observation value using $\hat{{\bm{y}}}=X{\bm{b}}$.
However, this approximated structure was formally the same as the mean structure of GCM. See Von Rosen [24] for details on the model. In Section 5, we compared the parameter matrices between NMF and GCM. In relation to this, GCM can also be considered as a varying coefficient model, as proposed by Hastie and Tibshirani [10], because the coefficient vector varies depending on the covariates. For details on the varying coefficients of GCM, see Satoh and Yanagihara [22]. Satoh and Tonda [21] described varying coefficients in an extended Growth Curve Model and Satoh et al [23] considered varying coefficients in the survival analysis.
An explanation has been added regarding the relationship with the case without covariates. If we set the covariate matrix as the identity matrix, that is, $R=N$ and $A=I_{N}$ then, it holds $B=\Theta$ and becomes equivalent to NMF without covariates. When we use covariates, the possible values of the coefficient vector are restricted. Therefore, the approximation may be worse than in cases without covariates. Therefore, instead of using available data as covariates, Satoh [20] proposed using a covariate matrix with a Gaussian kernel function that is useful for describing nonlinear structures. That is, By using individual data, ${\bm{u}}_{1},\ldots,{\bm{u}}_{N}$, we defined a Gaussian kernel for ${\bm{u}}$
$K({\bm{u}}_{n},{\bm{u}})=\exp(-\beta|{\bm{u}}_{n}-{\bm{u}}|^{2}),\qquad n=1,\ldots,N,$ (4)
where $\beta>0$ and give the covariate for individual $n$ as
${\bm{a}}_{n}=\displaystyle\left\{K({\bm{u}}_{1},{\bm{u}}_{n}),\ldots,K({\bm{u}}_{N},{\bm{u}}_{n})\right\}^{\prime}.$ (5)
Note that even if we use the same number of covariates as the number of individuals, we can obtain the optimal parameter matrix using the update formula in Section 3. A multiple regression model with a similar covariate matrix using ridge estimators is known as the Kernel Ridge Regression Model (see Murphy [17]. Additionally, Chen et al [3] conducted a study using a general kernel-based method in NMF without covariates.
Thus, we obtain the covariate matrix using Gaussian kernel functions $A(\beta)=({\bm{a}}_{1},\ldots,{\bm{a}}_{N})$. Here, $R=N$ and $A^{\prime}=A$. When ${\bm{u}}_{1},\ldots,{\bm{u}}_{N}$ differs, $A(\beta)\approx I_{N}$ holds when $\beta$ is large. Hence, each individual has a different coefficient vector and no longer depends on covariates. Therefore, as $\beta$ increases, the approximation accuracy of the observed values increases; however, new covariates cannot be predicted. Therefore, in Section 5 and , we optimize $\beta$ using cross-validation to improve the prediction accuracy. In Section 5.4, A more complex covariate matrix is introduced and symmetric matrix $A$ using Gaussian kernel functions.
3 Optimization of base matrix and parameter matrix
We introduce update formulas for optimizing the basis matrix $X$ and parameter matrix $\Theta$ when the observation matrix $Y$ and the known covariate matrix $A$ are provided.
The prediction matrix of $Y$ using NMF is
$\hat{Y}=({\bm{\hat{y}}}_{1},\dots,{\bm{\hat{y}}}_{N})=(\hat{y}_{p,n})_{P\times N}=XB=X\Theta A$ (6)
We measure the discrepancy between $Y$ and $\hat{Y}$ by 1) the squared Euclidean distance and 2) Kullback-Leibler (KL) divergence. Minimizing each objective function is equivalent to applying the maximum likelihood method when the observation is distributed as 1) Normal, and 2) Poisson. The update formulas for tri-NMF with a squared Euclidean distance were also presented by Ding et al [6] and Čopar et al [5] when $X$ and $A$ are orthogonal matrices. For the numerical computation of NMF without covariates, the package NMF of the statistical software R (R Core Team [19]) by Gaujoux and Seoighe [8] is available.
In addition to the above objective functions, the Itakura-Saito pseudo-distance (see Févotte et al [7]), which corresponds to the maximum likelihood method of the exponential distribution, and the $\beta$-divergence (see Cichocki and Amari [4], and Tweedie distributions by Abe and Yadohisa [1], are all additionally used in NMF.
3.1 Squared Euclidean distance
We consider the minimization of the squared Euclidean distance $\mbox{tr}(Y-\hat{Y})^{\prime}(Y-\hat{Y})$ between the observation matrix $Y$ and approximation matrix $\hat{Y}$ with an $L^{2}$ regularization term:
$D_{EU}(Y,\hat{Y})=\mbox{tr}(Y-\hat{Y})^{\prime}(Y-\hat{Y})+\gamma\mbox{tr}\Theta^{\prime}\Theta,$ (7)
where $\gamma$ is a non-negative penalty parameter, and $\Theta^{\prime}$ denotes the transposed matrix of $\Theta$.
When the coefficient matrices $B=\Theta A$ and $\hat{Y}$ are given, the update formula for the basis matrix $X$ remains the same as that without using covariates.
$x_{p,q}\longleftarrow x_{p,q}\displaystyle\frac{\displaystyle\sum_{n}y_{p,n}b_{q,n}}{\displaystyle\sum_{n}\hat{y}_{p,n}b_{q,n}},$ (8)
[…]
$X\longleftarrow X\odot(YB^{\prime}\oslash\hat{Y}B^{\prime}).$ (9)
Next, we present the update formula for the parameter matrix $\Theta$ when the basis matrices $X$ and $\hat{Y}$ are as follows:
$\theta_{q,r}\longleftarrow\displaystyle\theta_{q,r}\frac{\displaystyle\sum_{p,n}y_{p,n}x_{p,q}a_{r,n}}{\displaystyle\sum_{p,n}\hat{y}_{p,n}x_{p,q}a_{r,n}+\gamma\theta_{q,r}},$ (10)
where $\sum_{p,n}$ denotes $\sum_{p}\sum_{n}$. Hence, it can be written in matrix form as follows:
$\Theta\longleftarrow\Theta\odot\{(X^{\prime}YA^{\prime})\oslash(X^{\prime}\hat{Y}A^{\prime}+\gamma\Theta)\}$ (11)
3.2 Kullback-Leibler (KL)-type divergence
Given the KL-type divergence with $L^{2}$ regularization term between the observation matrix $Y$ and the approximation matrix $\hat{Y}$:
$D_{KL}(Y,\hat{Y})=\displaystyle\sum_{p,n}\left(y_{p,n}\log\frac{y_{p,n}}{\hat{y}_{p,n}}-y_{p,n}+\hat{y}_{p,n}\right)+\gamma\sum_{q,r}\theta_{q,r}^{2}.$ (12)
We aim to minimize it. First, when the coefficient matrices $B=\Theta A$ and $\hat{Y}$ are provided, the update formula for the basis matrix $X$ remains the same as that without covariates.
$x_{p,q}\longleftarrow x_{p,q}\displaystyle\frac{\displaystyle\sum_{n}b_{q,n}y_{p,n}/\hat{y}_{p,n}}{\displaystyle\sum_{n}b_{q,n}},$ (13)
or equivalently, $X\longleftarrow X\odot\{(Y\oslash\hat{Y})B^{\prime}\}\oslash\{{\bm{1}}_{P}{\bm{s}}_{r}(B)^{\prime}\}$. Here, ${\bm{1}}_{P}$ is a vertical vector of length $P$ with all components equal to 1. ${\bm{s}}_{r}(\cdot)$ is a function that returns the row sum of a matrix as a vertical vector. Therefore, ${\bm{1}}_{P}{\bm{s}}_{r}(B)^{\prime}$ is a $P\times Q$ matrix, where each row corresponds to the horizontal vector of the row sum of $B$.
Next, an update formula is introduced for the parameter matrix $\Theta$ when the basis matrices $X$ and $\hat{Y}$ are given.
$\theta_{q,r}\longleftarrow\theta_{q,r}\displaystyle\frac{\displaystyle\sum_{p,n}x_{p,q}a_{r,n}y_{p,n}/\hat{y}_{p,n}}{\displaystyle\sum_{p,n}x_{p,q}a_{r,n}+2\gamma\theta_{q,r}},$ (14)
or equivalently, $\Theta\longleftarrow\Theta\odot\{X^{\prime}(Y\oslash\hat{Y})A^{\prime}\}\oslash\{{\bm{s}}_{c}(X)^{\prime}{\bm{s}}_{r}(A)^{\prime}+2\gamma\Theta\}$ where ${\bm{s}}_{c}(\cdot)$ is a function that returns the column sum of the matrix as a horizontal vector.
4 Comparison with the Growth Curve Model
The Growth Curve Model (GCM) proposed by Potthoff and Roy [18] is as follows:
$Y\sim N_{P\times N}(X\Theta A,\Sigma\otimes I_{N}).$ (15)
The mean structure of GCM and the approximation by NMF with covariates are formally identical and can be described as $X\Theta A$. A comparison of NMF with the covariates and GCM is summarized in Table 1. In GCM, $X$ is called the within-individual covariate matrix, and A is the between-individual covariate matrix, both of which are known and given by the analyst. However, $X$ in NMF is unknown and optimized. Additionally, for uniqueness, each column vectorts of X is restricted to a sum or norm of 1, as explained in Section 5.
Table 1: Comparison of NMF with covariates and GCM
NMF with covariates GCM
Model $Y\approx X\Theta A$ $\mbox{E}(Y)=X\Theta A$
$Y$ non-negative values continuous values
$X$ restricted, unknown, optimized known
$\Theta$ unknown, optimized unknown, estimated
$A$ known known
[…]
X optimized by NMF can additionally be used in GCM and the model fitting of GCM can be improved. We demonstrate its usefulness using data analysis examples in Section 5.1. Further, the NMF and GCM are similar, $\Theta$ is unknown, and $A$ is known.
It is known that using the squared Euclidean distance as the objective function of NMF is equivalent to assuming a normal distribution, as in GCM, with the further difference that in GCM each column vector of $Y$ has a covariance matrix $\Sigma$. Note that $\Sigma$ is unstructured and contains $p(p+1)/2$ parameters.
The maximum likelihood estimators of $\Theta$ and $\Sigma$ on the GCM are obtained by
$\hat{\Theta}=(X^{\prime}S^{-1}X)^{-1}X^{\prime}S^{-1}YA^{\prime}(AA^{\prime})^{-1},\qquad\hat{\Sigma}=\frac{1}{N}(Y-X\hat{\Theta}A)(Y-X\hat{\Theta}A)^{\prime},$ (16)
where $S=Y\{I_{N}-A^{\prime}(AA^{\prime})^{-1}A\}Y^{\prime}/(N-R)$ and $N-R>0$. Thus, the following two points are the greatest advantages of GCM: 1) it does not require repetitive calculations to obtain estimators, and 2) tests and confidence intervals can be constructed based on the distribution of the estimators.
When the basis and covariate matrices of the GCM are non-negative, the non-negative parameter matrix can be optimized using the NMF update formula. From equation (11), this can be written as
$\Theta\longleftarrow\Theta\odot\{(X^{\prime}YA^{\prime})\oslash(X^{\prime}\hat{Y}A^{\prime})\},\qquad\hat{Y}\longleftarrow X\Theta A,$ (17)
Here, $\gamma=0$ was assumed as the penalty parameter.
Furthermore, the multiple regression model can be described as a special case of GCM. If the dimension of the individual observation data is one, the observation matrix is a transversal vector ${\bm{y}}^{\prime}$ of length $N$ and the basis matrix becomes scalar and can be written as $1$. Here, the parameter matrix is a transversal vector ${\bm{\theta}}^{\prime}$ of length $R$, which corresponds to the multiple regression model: ${\bm{y}}^{\prime}\approx{\bm{\theta}}^{\prime}A$ or ${\bm{y}}\approx A^{\prime}{\bm{\theta}}$. Then, the update equation can be written from equation (17) as
${\bm{\theta}}\longleftarrow{\bm{\theta}}\odot(A{\bm{y}}\oslash A\hat{{\bm{y}}}),\qquad\hat{{\bm{y}}}\longleftarrow A^{\prime}{\bm{\theta}}.$ (18)
The resulting regression coefficients are restricted to non-negative values, which makes them less accurate approximations than least-squares estimators; however, they might have the advantage of being easy to interpret.
5 Examples of data analysis
Before applying the proposed method to real data, we supplement the implementation details. First, when implementing NMF, it is sometimes necessary to impose constraints on the basis matrix to make the decomposition unique, such as making each column of the basis matrix a unit vector or making the column sum equal to 1. Here, we normalize the basis matrix such that the column sum is 1 because it makes the interpretation of the basis matrix and coefficient matrix easier. To set the column sum to 1, the following update equation can be inserted immediately after (8) or (13):
$x_{p,q}\longleftarrow\displaystyle\frac{x_{p,q}}{\displaystyle\sum_{p}x_{p,q}},\qquad q=1,\ldots,Q,\qquad\mbox{or,}\qquad X\longleftarrow X\oslash\{{\bm{1}}_{p}{\bm{s}}_{c}(X)\}.$ (19)
Because the sum of the components of each column of the basis matrix is 1, the elements can be interpreted as proportions or probabilities.
Because the coefficient matrix is also non-negative, it is possible to evaluate the contribution of the $q$th basis for individual $n$ as a ratio
$b_{q,n}/\sum_{q}b_{q,n},\qquad q=1,\ldots,Q,$ (20)
which can be used as the membership probability in soft clustering. Using terminology from the Topic Model in text analysis, the basis can be called a topic, and the probability can be called a topic probability.
NMF with or without covariates, has a large number of parameters to be optimized, and is a good fit to the data. To substantially reduce the number of parameters, it is effective to 1) reduce the number of bases $Q$ which is known as rank selection, 2) reduce $\beta$ in the Gaussian kernel function, and 3) increase the penalty parameter $\gamma$ of the objective function. Rank selection is a major problem, and probabilistic models (see, Ito et al [13]) have been proposed. However, in the example presented here, Q=1,2 or 3 is assumed for ease of interpretation. Optimization of $\beta$ was performed using a cross-validation method, described in detail in Section 5.2. In all the examples presented here, $\gamma$ is set to 0, because adjusting $\beta$ is sufficient to suppress overfitting. In fact, as $\gamma$ increases, the parameter matrix $\Theta$ approaches 0 from the update equations in (10) and (14); consequently, $\hat{Y}$ also approaches 0, resulting in generally poor prediction accuracy.
Finally, in Section 5, the squared Euclidean distance is used as the objective function; however, the results do not change significantly when the KL-type divergence is used. Conversely, we showed the coefficients of determination, which are easier to interpret as absolute values than the squared Euclidean distance.
Four datasets are treated here, but ordinary GCM was only applied to the first 4.1. For the other three, the maximum likelihood estimator could not be obtained in GCM: GCM is suitable for making statistical inferences, but was difficult to apply to observations of various sizes.
5.1 Growth curve data on an orthodontic measurement
The dataset was originally introduced in Table 1 of Potthoff and Roy [18] and is available as Orthodont in the nmle package in R. According to its description, Orthodont is a dataset collected at the University of North Carolina Dental School. There were 27 children (16 males and 11 females) aged 8-–14 years. Every two years, they measured the distance between the pituitary and pterygomaxillary fissures. Thus, the observation matrix $Y$ where $P=4$ and $N=27$ consists of the distance among 16 males and 11 females at the ages of 8, 10, 12, and 14. Script 1 shows the R scripts used for data preparation. An observation matrix $Y$ was created from the Orthodont data, and four individuals were selected, two each from the male and female data. Although the observed values were expected to increase monotonically with age, they were measured using X-rays, which introduced an error in the measurement. For example, M01 represents the first male individual; however, the observed value at 10 years of age is 25, which is shorter than the observed value of 26 at 8 years of age.
Script 1: Creation of observation matrix
> library(nlme); Y <- matrix(Orthodont\$distance,nrow=4)
> rownames(Y) <- unique(Orthodont\$age)
> colnames(Y) <- unique(Orthodont\$Subject)
> index <- c(1,16,17,27); Y[,index]
M01 M16 F01 F11
8 26 22.0 21.0 24.5
10 25 21.5 20.0 25.0
12 29 23.5 21.5 28.0
14 31 25.0 23.0 28.0
Next, we use the nmfkc package, which we have developed and published on GitHub [^1] https://github.com/ksatohds/nmfkc, to perform ordinary NMF without covariates given by (1). In other words, $X$ and $B$ approximate to $Y$ in Script 2. First, install the package available from GitHub and use $Y$ given in Script 1 as the argument of the nmfkc function. The number of bases was set to $Q=2$, which is equal to that of a linear curve. The optimized basis and coefficient matrices are then presented. Note that the sum of each column of the basis matrix is 1.
Script 2: Installation of the required packages and execution of NMF without covariates
> #install.packages("remotes")
> #remotes::install_github("ksatohds/nmfkc")
> library(nmfkc); result <- nmfkc(Y,Q=2,epsilon=1e-8)
> result\$r.squared
[1] 0.9064937
> round(result\$X,3)
1 2
8 0.302 0.159
10 0.215 0.267
12 0.266 0.248
14 0.217 0.327
> round(result\$B[,index],1)
M01 M16 F01 F11
1 57.4 50.1 49.3 56.6
2 53.7 41.9 36.2 48.9
Fig.1 shows the fit of the NMF without covariates. The resulting coefficient of determination for the corresponding components of $Y$ and $\hat{Y}=XB$ was $0.906$, which indicates excellent approximation accuracy.
Figure 1: Observation and fitted curves for each individual based on NMF without covariates. Circles indicate observed values and solid lines indicate fitted curves.
Y used in the NMF and the optimized X and B. Because the coefficients $57.4$ and $53.7$ are also non-negative, their proportions, $0.517$ and $0.483$, can be used as probabilities for soft clustering.
$Y=\begin{pmatrix}26\cdots 22.0&21.0\cdots 24.5\\ 25\cdots 21.5&20.0\cdots 25.0\\ 29\cdots 23.5&21.5\cdots 28.0\\ 31\cdots 25.0&23.0\cdots 28.0\\ \end{pmatrix},X=\begin{pmatrix}0.302&0.159\\ 0.215&0.267\\ 0.266&0.248\\ 0.217&0.327\\ \end{pmatrix},B=\begin{pmatrix}57.4\cdots 50.1&49.3&\cdots 56.6\\ 53.7\cdots 41.9&36.2&\cdots 48.9\\ \end{pmatrix}.$
Script 3 performs NMF using a covariate matrix composed of male dummy variables. Note that the optimized parameter matrix $\Theta$ is taken as C.
Script 3: NMF with covariates
> A <- rbind(rep(1,ncol(Y)),substr(colnames(Y),1,1)=="M")
> colnames(A) <- colnames(Y)
> rownames(A) <- c("(Intercept)","Male")
> A[,index]
M01 M16 F01 F11
(Intercept) 1 1 1 1
Male 1 1 0 0
> result <- nmfkc(Y,A,Q=2,epsilon=1e-8)
> result\$r.squared
[1] 0.4267753
> round(result\$X,3)
Basis1 Basis2
8 0.175 0.284
10 0.160 0.317
12 0.287 0.227
14 0.377 0.172
> round(result\$C,3)
(Intercept) Male
Basis1 41.550 8.627
Basis2 49.036 0.665
The optimized X and parameter matrices and the covariate matrix consisting of dummy variables are described as follows:
$X=\begin{pmatrix}0.175&0.284\\ 0.160&0.317\\ 0.287&0.227\\ 0.377&0.172\\ \end{pmatrix},\Theta=\begin{pmatrix}41.550&8.627\\ 49.036&0.665\\ \end{pmatrix},A=\begin{pmatrix}1\cdots 1&1\cdots 1\\ 1\cdots 1&0\cdots 0\\ \end{pmatrix}.$
The fitted curves are shown in Fig.2. There was no difference in the slope of the curves for the sexes aged 8–10 years; however, after the age of 10 years, the slope for males was slightly higher.
From (16), the maximum likelihood estimator, $\hat{\Theta}$, is obtained using $X$ optimized by NMF for GCM,
$\hat{\Theta}=\begin{pmatrix}41.522&8.672\\ 49.066&0.616\\ \end{pmatrix}.$
The optimized $\Theta$ values obtained from the NMF and GCM estimates $\hat{\Theta}$ were almost identical. The coefficient of determination based on $X$ optimized by NMF is 0.427, which decreases to 0.422 when a line with the same base number 2 is assumed for $X$. Therefore, even when using a growth curve model, it may be better to use X optimized by NMF than X given by the analyst.
Figure 2: Observation and fitted lines by NMF with covariates. Gray lines indicate individual observations and bold lines show the fitted curves.
5.2 Daily temperature in Canada
NMF was applied to daily mean temperature data from 1960 to 1994 at 35 weather stations in Canada. The data are included in the CanadianWeather of the fda package, and they also contain the location information of the weather stations.
In Script 4, we define an observation matrix $Y$ with $P=365$ and $N=35$. The minimum temperature value was $-34.8$, which was negative; therefore, it was subtracted from the temperature such that the minimum value of Y was 0. NMF was applied with a base number $Q=2$ and the coefficient of determination was $0.985$, which is very high. Fig.3 shows the observed values and fitted curves at the 35 weather stations. The horizontal axis of each figure is aligned to 365 days from January 1 to December 31, and the vertical axis is also aligned to 25 degrees for the maximum value and -35 degrees for the minimum value. The horizontal dotted line indicates 0 degrees Celsius. Although the observed and fitted values almost overlap, a difference can be seen at the 35th resolution.
Script 4: NMF without covariates when dimension $P$ is high
> library(fda); data(CanadianWeather)
> d <- CanadianWeather\$dailyAv[,,1]; Y <- d-min(d)
> library(nmfkc); result <- nmfkc(Y,Q=2); result\$r.squared
[1] 0.9853628
Figure 3: Observed and fitted values for 35 weather stations in Canada. The observed values are shown in gray and the fitted curve by NMF is shown in black. The dotted line indicates 0 degrees.
Fig.4 shows (a) the column vectors of the basis matrix, and (b) the probabilities at the weather stations. The sum of each basis vector is 1. The fitted value for each weather station was approximated as the sum of the coefficient multiples of the two bases. The two coefficients are non-negative and can be interpreted as probabilities, as shown in the star chart with normalized longitudes and latitudes. As shown in Fig.4(a), basis 1 represents the temperature variation in the Northern Hemisphere, with lower temperatures in winter and higher temperatures in summer. Basis 2, conversely, is shaped to push up winter temperatures. In Fig.4(b), the probability of basis 1 is higher north of the interior, and the probability of basis 2 increases southward. Additionally, the probability of basis 2 tends to be higher on the west coast than on the east coast, even in the same southern area. These trends suggest that the coefficient matrix can be explained by location information: longitude and latitude.
Figure 4: (a) Two bases described as column vectors of $X$ and (b) probabilities obtained as the ratio of the coefficients of two bases at weather stations.
Script 5 performs NMF with the location information based on the trends identified in Fig.4(b). First, the longitudes and latitudes were normalized such that the minimum and maximum values were 0 and 1, respectively. Let $U$ be a matrix where the first row is the normalized longitude, and the second row is the normalized latitude. Then, the covariate matrix $A$ was created from the normalized matrix $U$ using a Gaussian kernel function with the parameter $\beta=6.1$. Varying the parameter $\beta$ can significantly change the fit to the data. Fig.5(a) shows that increasing $\beta$ decreases the value of the objective function and eventually converges to the value of the NMF without covariates. This can be understood from the fact that the covariate matrix $A$ created from the Gaussian kernel function converges to the identity matrix. Contrastingly, Fig.5(b) minimizes the objective function using the cross-validation method by varying $\beta$. The objective function reaches its minimum value when $\beta=6.1$. Cross-validation was performed using the nmfkc.cv function.
Script 5: NMF with covariates given by the Gaussian kernel function
> u <- CanadianWeather\$coordinates[,2:1]
> u[,1] <- -u[,1]; r <- apply(u,2,range)
> U <- t(scale(u,center=r[1,],scale=r[2,]-r[1,]))
> A <- nmfkc.kernel(U,beta=6.1)
> library(nmfkc); result <- nmfkc(Y,A,Q=2); result\$r.squared
[1] 0.9340437
Figure 5: (a) Objective function values and (b) 10-fold cross-validation which vary with $\beta$ in the Gaussian kernel function. The dotted line in (a) shows the value for NMF without covariates.
From equation (3), the probability of basis 2 can be derived on new locations and it is obtained in B.prob of Script 6 and those can be visualized as Fig.6. The general trend is the same as in Fig.4(b).
Script 6: Soft clustering on new covariates
> v <- seq(from=0,to=1,length=20)
> V <- t(cbind(expand.grid(v,v)))
> A <- nmfkc.kernel(U,V,beta=best.beta);
> B <- result\$C %*A; B.prob <- prop.table(B,2)
Figure 6: Geographic distribution of the probability of basis 2. The contours lines shows the probability of basis 2 and those are obtained as smooth curves based on Gaussian kernel functions.
Additional information is provided on NMF with covariates. Although the covariate matrix $A$ was created from $U$ using the Gaussian kernel function, $U$ could be used directly as the covariate matrix, and the coefficient of determination in this case was not high (0.715).
5.3 US presidential inaugural address texts
We treated the inaugural speeches of U.S. presidents as textual data observed over time. The text data were stored as data_corpus_inaugural in the quanteda package and contained 59 speeches from President Washington in 1789 to President Biden in 2021. In Script 7, we first extracted from the text data 40 frequent words with 5 or more characters and 100 or more occurrences. Generally, a frequency table in which the rows are words, and the columns are documents is called a term–document matrix. If we apply NMF to the term document matrix, then such can be interpreted as a Topic Model. In other words, each column vector of the basis matrix is the probability of a word constituting a topic, and the proportion of coefficients can be interpreted as topic probability. The rows of the basis matrix are all probabilities of the same word for each topic, and the proportion of probabilities provides the magnitude of the word’s contribution to each topic.
Script 7: NMF as Topic Model
> library(quanteda)
> tok <- tokens(corpus(data_corpus_inaugural))
> tok <- tokens_remove(tok,
+ pattern=stopwords("en",source="snowball"))
> df <- dfm_select(dfm(tok),min_nchar=5)
> df <- dfm_trim(df,min_termfreq=100)
> Y <- t(df[,order(colSums(df),decreasing=T)])
> Y[1:5,c(1,2,59)]
1789-Washington 1793-Washington 2021-Biden
people 4 1 9
government 8 1 0
great 3 0 4
states 2 0 1
world 1 0 8
> library(nmfkc); result <- nmfkc(Y,Q=3); result\$r.squared
[1] 0.6563144
> U <- matrix(data_corpus_inaugural\$Year,nrow=1)
> A <- nmfkc.kernel(U,beta=0.00015)
> result <- nmfkc(Y,A,Q=3); result\$r.squared
[1] 0.3280116
[…]
NMF was then applied by creating a covariate matrix $A$ by using the Gaussian kernel function from the year of inauguration. The coefficient of determination was 0.328, which was worse. Bar chart of topic probabilities for each president obtained under the model are shown in Fig.7(b). The general trend is similar to that shown in Fig.7(a), but the Gaussian kernel function smoothes out the vertical variation seen in Fig.7(a), making Fig.7(b) more straightforward to interpret.
Figure 7: Bar chart of topic probability for each president by NMF (a) with and (b) without covariates.
5.4 Bone mineral density in adolescents
Relative changes in the spinal bone mineral density of 261 North American adolescents (116 males and 145 females) were used as observations. The bone mineral content is a measure of the amount of minerals contained in each amount of bone and is used to diagnose osteoporosis. The observed value was the relative change in bone mineral content between two consecutive measurements divided by the average age. The number of observation points was 107 for 1 measurement, 84 for 2 measurements, and 70 for 3 measurements. Therefore, the total number of observations was 485. The data are presented as an example of the application of smoothing in Chapter 5 of Hastie et al [11]. The data were published as bone in the ElemStatLearn package.  [^2] https://cran.r-project.org/src/contrib/Archive/ElemStatLearn/.
In Script 8, the dimension of the observation matrix $Y$ is $P=1$, resulting in a scalar basis matrix with $X=1$ and $Q=1$. Hence, the NMF with covariates is equivalent to a multiple regression model that can be written as the update equation (18). The Gaussian kernel function for age and the function of the product of age and male dummy variables were used to obtain smoothed curves for gender, see Fig.8.
Script 8: NMF with the covariates as Multiple Regression Model
> library(ElemStatLearn)
> d <- bone; d\$male <- ifelse(d\$gender=="male",1,0)
> d <- d[order(d\$male,d\$age),]
> Y <- matrix(d\$spnbmd-min(d\$spnbmd),nrow=1)
> library(nmfkc)
> A1 <- nmfkc.kernel(matrix(d\$age,nrow=1),beta=0.32)
> A2 <- nmfkc.kernel(matrix(d\$age*d\$male,nrow=1),beta=0.32)
> result <- nmfkc(Y,A=rbind(A1,A2),Q=1); result\$r.squared
[1] 0.4264902
Figure 8: Gender smoothing curves for relative change in spinal bone mineral density. The filled circles indicate the observed values, and the solid lines indicate the fitted curves. Black indicates males and gray indicates females, and the dotted line indicates no relative change.
6 Conclusion
In this study, NMF with covariates was applied to the longitudinal measurement and text data, citing similarities with the growth curve model. Although the growth curve model has many advantages in statistical inference owing to its explicit maximum likelihood estimator, the update formula of NMF allows for stable optimization despite the large number of parameters involved. It is also possible to use covariates that exceed the amount of data, as is the case with Gaussian kernel functions. We have also developed an R package and published it in GitHub to implement these methods and have shown through data examples that they can be easily implemented. NMF is useful in exploratory data analysis because it has aspects of both the regression model and clustering method. NMF with covariates can also make predictions using covariates.
[…]
Acknowledgements
This work was partly supported by JSPS KAKENHI Grant Numbers JP22K11930 and the project research fund by the Research Center for Sustainability and Environment at Shiga University.